# A Theoretical Framework for Data Efficient Multi-Source Transfer Learning Based on Cramér-Rao Bound

### v2

@@ Abstract
Multi-source transfer learning provides an effective solution to data scarcity in real-world supervised learning scenarios by leveraging multiple source tasks. In this field, existing works typically use all available samples from sources in training, which constrains their training efficiency and may lead to suboptimal results. To address this, we propose a theoretical framework that answers the question: what is the optimal quantity of source samples needed from each source task to jointly train the target model? Specifically, we introduce a generalization error measure that aligns with cross-entropy loss, and minimize it based on the Cramér-Rao Bound to determine the optimal transfer quantity for each source task. Additionally, we develop an architecture-agnostic and data-efficient algorithm OTQMS to implement our theoretical results for training deep multi-source transfer learning models. Experimental studies on diverse architectures and two real-world benchmark datasets show that our proposed algorithm significantly outperforms state-of-the-art approaches in both accuracy and data efficiency. The code and supplementary materials are available in https://anonymous.4open.science/r/Materials.
Machine Learning, ICML
@@ 1 Introduction
Nowadays, various machine learning algorithms have achieved remarkable success by leveraging large-scale labeled training data. However, in many practical scenarios, the limited availability of labeled data presents a significant challenge, where transfer learning emerges as an effective solution (Zhao et al., 2020). Transfer learning aims to leverage knowledge from tasks with abundant data or being well-trained, known as the source tasks, to improve the performance of a new learning task, known as the target task. Given its numerous applications, transfer learning has gained wide popularity and seen success in a variety of fields, such as computer vision (Wang & Deng, 2018), natural language processing (Sung et al., 2022), recommendation system (Fu et al., 2024) and anomaly detection (Vincent et al., 2020). Traditionally, transfer learning has focused on the transfer between a single source task and a target task. However, there is a growing emphasis on multi-source transfer learning, which leverages multiple source tasks to enhance the training of the target task (Sun et al., 2015).
@@
Figure 1: More source samples does not always mean better performance. Incorporating all source samples may bring negative impact, which is illustrated by the comparison of two strategies, using target with all source samples (blue) and using target only (red), evaluated on the equally divided 5-task CIFAR10 dataset. Theoretically, although incorporating more source samples reduces model variance by expanding the training data, the discrepancy between the source and target tasks introduces additional bias.
@@
MCW (Lee et al., 2019) Supervised
Leep (Nguyen et al., 2020) Supervised
Tong (Tong et al., 2021) Supervised
DATE (Han et al., 2023) Unsupervised
H-ensemble (Wu et al., 2024) Supervised
@@
In multi-source transfer learning, traditional methods usually jointly train the target model using all available samples from sources without selection (Zhang et al., 2024; Shui et al., 2021; Li et al., 2021b). This evidently poses a severe limitation to training efficiency, considering the vast number of available samples from various potential source tasks in real-world scenarios (Peng et al., 2019). Moreover, directly assuming the use of all available samples seriously constrains their solution space, which possibly leads to suboptimal results as illustrated in Figure 1. Therefore, it is critical to establish a theoretical framework to answer the question: what is the optimal transfer quantity of source samples from each source task needed in training the target task?
In this work, we establish a theoretical framework for the multi-source transfer learning problem. We formulate the transfer quantity of each source task as an optimization variable and propose a method to determine its optimal value. Specifically, we introduce the expectation of Kullback-Leibler (K-L) divergence between the distribution learned from training samples and the true distribution of target task samples as a measure of generalization error. This measure is then minimized based on the Cramér-Rao Bound to derive the optimal transfer quantity of each source task. Building on this, we propose a practical algorithm, named Optimal Transfer Quantities for Multi-Source Transfer learning (OTQMS), to implement our theoretical results for training deep multi-source transfer learning models. Notably, OTQMS is data-efficient and compatible with various model architectures, including Vision Transformer (ViT) and Low-Rank Adaptation (LoRA). It also demonstrates advantages in task generality and shot generality (as illustrated in Table 1), since we establish our theoretical framework without restricting it to a specific target task type or limiting the range of target sample quantity.
In summary, our main contributions are as follows:
  • a)
    A novel theoretical framework that optimizes transfer quantities is introduced for multi-source transfer learning. We propose a new K-L divergence measure of the generalization error. By minimizing it, we develop a method to determine the optimal transfer quantity of each source task.
  • b)
    Based on the framework, we propose OTQMS, an architecture-agnostic and data-efficient algorithm for target model training in multi-source transfer learning.
  • c)
    Experimental studies on two real-world datasets and various model architectures demonstrate that OTQMS achieves significant improvements in both accuracy and data efficiency. In terms of accuracy, OTQMS outperforms state-of-the-art by an average of $1.5\%$ on DomainNet and $1.0\%$ on Office-Home. In terms of data efficiency, OTQMS reduces the average training time by $35.19\%$ and the average sample usage by $47.85\%$ on DomainNet. Furthermore, extensive supplementary experiments demonstrate that our theory can facilitate both full model training and parameter-efficient training.
@@ 2 Related Work
2.1 Transfer Learning Theory
Existing theoretical works can be categorized into two groups. The first group focuses on proposing measures to quantify the similarity between the target and source tasks. Within this group, some measures have been introduced, including $l_{2}$-distance (Long et al., 2014), optimal transport cost (Courty et al., 2016), LEEP (Nguyen et al., 2020), Wasserstein distance (Shui et al., 2021), and maximal correlations (Lee et al., 2019). This work belongs to the second group focusing on developing new generalization error measures. Within this group, the measures having been introduced include $f$-divergence (Harremoës & Vajda, 2011), mutual information  (Bu et al., 2020), ${\mathcal{X}}^{2}$-divergence (Tong et al., 2021), $\mathcal{H}$-score (Bao et al., 2019; Wu et al., 2024). However, the potential of K-L divergence as a generalization error measure has not been sufficiently explored.
2.2 Multi-source Transfer Learning
Classified by the object of transfer, existing multi-source transfer learning methods mainly focus on two types: model transfer vs sample transfer (Zhuang et al., 2020). Model transfer assumes there is one or more pre-trained models on the source tasks and transfers their parameters to the target task via fine-tuning (Wan et al., 2022). This work focuses on the latter, which is based on joint training of the source task samples with those of the target task (Zhang et al., 2024; Shui et al., 2021; Li et al., 2021b). Classified by the strategy of transfer, existing methods mainly focus on two types: alignment strategy and matching-based strategy (Zhao et al., 2024). Alignment strategy aims to reduce the domain shift among source and target domains (Li et al., 2021a; Zhao et al., 2021; Li et al., 2021b). This work is more similar to the latter, focusing on determining which source domains or samples should be selected or assigned higher weights for transfer (Guo et al., 2020; Shui et al., 2021; Tong et al., 2021; Wu et al., 2024). However, most existing works either utilize all samples from all sources or perform task-level selection, whereas this work explores a framework that optimizes the transfer quantity of each source task.
@@ 3 Problem Formulation
Consider the transfer learning setting with one target task ${\mathcal{T}}$, and $K$ source tasks $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$. The target task $\mathcal{T}$ is not restricted to a specific downstream task category. Generally, we formulate it as a parameter estimation problem under a distribution model $P_{X;{\underline{\theta}}}$. For example, when ${\mathcal{T}}$ is a supervised classification task, $P_{X;{\underline{\theta}}}$ corresponds to the joint distribution model of input features $Z$ and output labels $Y$, i.e., $X=(Z,Y)$. Our objective is to estimate the true value of $\underline{\theta}$, which corresponds to optimizing the neural network parameters for target task ${\mathcal{T}}$. Here, $\theta$ denotes 1-dimensional parameter, and $\underline{\theta}$ denotes high dimensional parameter.
Furthermore, we assume that the source tasks and the target task follow the same parametric model and share the same input space $\mathcal{X}$. Without loss of generality, we assume $\mathcal{X}$ to be discrete, though our results can be readily extended to continuous spaces. The target task ${\mathcal{T}}$ has $N_{0}$ training samples $X^{N_{0}}=\{x_{1},\dots,x_{N_{0}}\}$ i.i.d. generated from some underlying joint distribution $P_{X;{\underline{\theta}}_{0}}$, where the parameter ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$. Similarly, the source task $\mathcal{S}_{k}$ has $N_{k}$ training samples $X^{N_{k}}=\{x^{k}_{1},\dots,x^{k}_{N_{k}}\}$ i.i.d. generated from some underlying joint distribution $P_{X;{\underline{\theta}}_{k}}$, where $k\in[1,K]$, and the parameter ${\underline{\theta}}_{k}\in\mathbb{R}^{d}$. In this work, we use the Maximum Likelihood Estimator (MLE) to estimate the true target task parameter $\theta_{0}$. Moreover, the following lemma demonstrates that, for large sample size, the MLE can achieve the theoretical lower bound of the Mean Squared Error (MSE) asymptotically, known as the Cramér-Rao Bound.
Lemma 3.1.
(Cover & Thomas, 2006) Let $\hat{{\underline{\theta}}}$ be an unbiased estimator of ${\underline{\theta}}_{0}$ based on $n$ i.i.d. samples. Then, the MSE matrix of $\hat{{\underline{\theta}}}$ satisfies the following lower bound:
$\text{MSE}(\hat{{\underline{\theta}}})^{d\times d}\triangleq\mathbb{E}\left[\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)\left(\hat{{\underline{\theta}}}-{\underline{\theta}}_{0}\right)^{T}\right]\geq\frac{1}{n}J({\underline{\theta}}_{0})^{-1},$ (1)
where $\left(\hat{{\underline{\theta}}}-{{\underline{\theta}}_{0}}\right)$ is $d-dimensional$ vector, and the matrix inequality $A\geq B$ means that $A-B$ is positive semi-definite. In addition, the $J({\underline{\theta}})$ is fisher information matrix defined as:
@@
$J({\underline{\theta}})^{d\times d}$ $=\mathbb{E}\left[\left(\frac{\partial}{\partial{\underline{\theta}}}\log P(X;{\underline{\theta}})\right)\left(\frac{\partial}{\partial{\underline{\theta}}}\log P(X;{\underline{\theta}})\right)^{T}\right].$ (2)
Moreover, as the sample size $n\to\infty$, the MLE $\hat{{\underline{\theta}}}$ can achieve this bound asymptotically, i.e.,
$\text{MSE}(\hat{{\underline{\theta}}})=\frac{1}{n}J({\underline{\theta}}_{0})^{-1}+o\left(\frac{1}{n}\right).$ (3)
It should be noted that in some works, the MSE is defined as a scalar, which corresponds to the trace of the MSE matrix and can also be bounded using this lemma.
@@
When we transfer $n_{1},\dots,n_{K}$ samples from the $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$, where $n_{k}\in\left[0,N_{k}\right]$, we denote these training sample sequences as $X^{n_{1}},X^{n_{2}},\dots,X^{n_{k}}$ Then, the MLE for multi-source transfer learning, in the simplest form, is a summation of the likelihood of all source tasks and target tasks, i.e.,
$\hat{{\underline{\theta}}}=\operatorname*{arg\,max}_{{\underline{\theta}}}$ ${\sum_{x\in X^{N_{0}}}}\log P_{X;{\underline{\theta}}}(x)$
$+{\sum_{k=1}^{K}\sum\limits_{x\in X^{n_{k}}}}\log P_{X;{\underline{\theta}}}(x).$ (4)
@@
In this work, our goal is to derive the optimal transfer quantities $n_{1}^{*},\dots,n_{K}^{*}$ of source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ to minimize certain divergence measure $\mathcal{L}$ between the distribution learned from training samples $P_{X;\hat{{\underline{\theta}}}}$ and the true distribution of target task $P_{X;{\underline{\theta}}_{0}}$, i.e.,
$n_{1}^{*},\dots,n_{k}^{*}=\operatorname*{arg\,min}_{n_{1},\dots,n_{k}}\mathcal{L}(P_{X;\hat{{\underline{\theta}}}},P_{X;{\underline{\theta}}_{0}}).$ (5)
@@ 4 Theoretical Analysis and Algorithm
In this section, we will first introduce a new K-L divergence measure for the optimization problem in (5). Then, we will analyze it based on the Cramér-Rao bound to derive the optimal transfer quantities for both single-source and multi-source scenarios. Finally, we will develop a practical algorithm based on the theoretical framework.
Definition 4.1.
(The K-L divergence) The K-L divergence measures the difference between two probability distributions $P(X)$ and $Q(X)$ over the same probability space. It is defined as:
@@
In this work, we apply the K-L divergence between the distribution model parameterized by MLE (3) and the true distribution model of target task to measure the generalization error, which directly corresponds to the testing error using cross-entropy loss (Cover & Thomas, 2006).
$D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)=-{\underbrace{H(P_{X;\hat{{\underline{\theta}}}})}_{\textnormal{entropy}}}+\underbrace{\mathbb{E}_{P_{X;\hat{{\underline{\theta}}}}}\left[P_{X;{\underline{\theta}}_{0}}\right]}_{\textnormal{cross-entropy}}.$
Finally, our generalization error measure is defined as follows.
$\mathcal{L}(P_{X;\hat{{\underline{\theta}}}},P_{X;{\underline{\theta}}_{0}})=\mathbb{E}\left[D\left(P_{X;\hat{{\underline{\theta}}}}\middle\|P_{X;{\underline{\theta}}_{0}}\right)\right].$ (6)
@@ 4.1 Single-Source Transfer Learning
The direct computation of the proposed K-L divergence measure is challenging. Fortunately, we can prove that the proposed K-L measure directly depends on the MSE in the asymptotic case. Moreover, the MSE has a widely used Cramér-Rao bound introduced in Lemma 3.1, which provides a bound on the estimator’s error, and is asymptotically tight. As our goal is to find a correspondence between transfer quantity and generalization error, for the rest of the paper, we mainly analyze the K-L measure based on the Cramér-Rao bound (3). To begin, we consider the setting with a target task ${\mathcal{T}}$ with $N_{0}$ samples generated from a model with 1-dimensional parameter.
Lemma 4.2.
(proved in Appendix B.1) Given a target task ${\mathcal{T}}$ with $N_{0}$ i.i.d. samples generated from 1-dimensional underlying model $P_{X;\theta_{0}}$, where $\theta_{0}\in\mathbb{R}$, and denoting $\hat{\theta}$ as the MLE (3) based on the $N_{0}$ samples, then the K-L measure (6) can be expressed as:
$\mathbb{E}\left[D\left(P_{X;\hat{\theta}}\middle\|P_{X;\theta_{0}}\right)\right]=\frac{1}{2N_{0}}+o\left(\frac{1}{N_{0}}\right).$ (7)
The result of Lemma 4.2 demonstrates that when there is only a target task, our K-L measure is inversely proportional to the number of training samples. Next, we consider the transfer learning scenario in which we have one target task ${\mathcal{T}}$ with $N_{0}$ training samples and one source task $\mathcal{S}_{1}$ with $N_{1}$ training samples. In this case, We aim to determine the optimal transfer quantity $n_{1}^{*}\in[1,N_{1}]$. To facilitate our mathematical derivations, we assume $N_{0}$ and $N_{1}$ are asymptotically comparable, and the distance between the parameters of target task and source task is sufficiently small (i.e., $|\theta_{0}-\theta_{1}|=O(\frac{1}{\sqrt{N_{0}}})$). Considering the similarity of low-level features among tasks of the same type, this assumption is made without loss of generality. Furthermore, as demonstrated in subsequent analysis, our conclusions remain valid even in extreme cases where the distance between parameters is large.
Theorem 4.3.
(proved in Appendix B.2) Given a target task ${\mathcal{T}}$ with $N_{0}$ i.i.d. samples generated from 1-dimensional underlying model $P_{X;\theta_{0}}$, and a source task $\mathcal{S}_{1}$ with $N_{1}$ i.i.d. samples generated from 1-dimensional underlying model $P_{X;\theta_{1}}$, where $\theta_{0},\theta_{1}\in\mathbb{R}$. $\hat{\theta}$ is denoted as the MLE (3) based on the $N_{0}$ samples from ${\mathcal{T}}$ and $n_{1}$ samples from $\mathcal{S}_{1}$, where $n_{1}\in[0,N_{1}]$, then the K-L measure (6) can be expressed as:
[…]
@@
From (8), we observe that the K-L measure decreases as $N_{0}$ increases, which aligns with our intuition that utilizing all available target samples is beneficial. In addition, the trend of (8) with respect to $n_{1}$ is more complex. We plot (8) as a function of $n_{1}$ under two different regimes, determined by the value of $N_{0}\cdot t$, as shown in Figure 2. Our goal is to explore the optimal value of $n_{1}^{*}$ to minimize (8).
Figure 2: The function curve figures of (8) under different regimes determined by the value of $N_{0}\cdot t$ (blue). The vertical axis denotes the value of K-L measure (8), while the horizontal axis denotes the variable $n_{1}$.
  • Case 1 ($N_{0}\cdot t\leq 0.5$): The K-L measure monotonically decreases as $n_{1}$ increases. Obviously, the optimal point is $n_{1}^{*}=N_{1}$. This indicates that when the source task and the target task are highly similar, i.e., when $t$ is small, an increase in the transfer quantity will positively impact the results.
  • Case 2 ($N_{0}\cdot t>0.5$): The K-L measure first decreases and then increases as $n_{1}$ increases. It attaining its minimum at $n_{1}^{\prime}=\frac{N_{0}}{2N_{0}t-1}$. It should be noted that when $t$ is large enough, the point $n_{1}^{\prime}$ approaches 00. This aligns with the intuition that when the discrepancy between the source and target tasks is substantial, avoiding transfer yields better results. Furthermore, if $n_{1}^{\prime}$ exceeds $N_{1}$, we should utilize all $N_{1}$ samples, so $n_{1}^{*}=\min\left(N_{1},\frac{N_{0}}{2N_{0}t-1}\right)$.
In Figure 3, we provide an intuitive explanation for the optimal $n_{1}^{*}$.
Figure 3: The three circles represent the errors corresponding to transfer quantity $n_{1}=0$, $n_{1}^{*}$, and $N_{1}$. The distance from each circle’s center to $\theta_{0}$ represents the bias of estimation, and the radius represents the variance. As the transfer quantity increases, the bias increases, while the variance decreases. The optimal $n_{1}^{*}$ achieves the best trade-off between them.
@@
Proposition 4.4.
(proved in Appendix B.3) In the case where the parameter dimension is d, i.e., ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$, with all other settings remaining the same as the Lemma 4.2, the K-L measure (6) can be expressed as
@@
Proposition 4.5.
(proved in Appendix B.4) In the case where the parameter dimension is d, i.e., ${\underline{\theta}}_{0},{\underline{\theta}}_{1}\in\mathbb{R}^{d}$, with all other settings remaining the same as the Theorem 4.3, the K-L measure (6) can be expressed as

### v3

# A High-Dimensional Statistical Method for Optimizing Transfer Quantities in Multi-Source Transfer Learning

@@ Abstract
Multi-source transfer learning provides an effective solution to data scarcity in real-world supervised learning scenarios by leveraging multiple source tasks. In this field, existing works typically use all available samples from sources in training, which constrains their training efficiency and may lead to suboptimal results. To address this, we propose a theoretical framework that answers the question: what is the optimal quantity of source samples needed from each source task to jointly train the target model? Specifically, we introduce a generalization error measure based on K-L divergence, and minimize it based on high-dimensional statistical analysis to determine the optimal transfer quantity for each source task. Additionally, we develop an architecture-agnostic and data-efficient algorithm OTQMS to implement our theoretical results for target model training in multi-source transfer learning. Experimental studies on diverse architectures and two real-world benchmark datasets show that our proposed algorithm significantly outperforms state-of-the-art approaches in both accuracy and data efficiency.
@@ 1 Introduction
Figure 1: More source samples does not always mean better performance. Incorporating all source samples may bring negative impact, which is illustrated by the comparison of two strategies, using target task samples with all source samples (blue) and using target task samples only (red), evaluated on the equally divided 5-task CIFAR10 dataset. Theoretically, although incorporating more source samples reduces model variance by expanding the training data, the discrepancy between the source and target tasks introduces additional bias.
Nowadays, various machine learning algorithms have achieved remarkable success by leveraging large-scale labeled training data. However, in many practical scenarios, the limited availability of labeled data presents a significant challenge, where transfer learning emerges as an effective solution [33]. Transfer learning aims to leverage knowledge from tasks with abundant data or being well-trained, known as the source tasks, to improve the performance of a new learning task, known as the target task. Given its numerous applications, transfer learning has gained wide popularity and seen success in a variety of fields, such as computer vision [28], natural language processing [24], recommendation systems [7] and anomaly detection [26]. Traditionally, transfer learning has focused on the transfer between a single source task and a target task. However, there is a growing emphasis on multi-source transfer learning, which leverages multiple source tasks to enhance the training of the target task [23].
In multi-source transfer learning, traditional methods usually jointly train the target model using all available samples from sources without selection [32, 22, 14]. This evidently poses a severe limitation to training efficiency, considering the vast number of available samples from various potential source tasks in real-world scenarios [19]. Moreover, directly assuming the use of all available samples seriously constrains their solution space, which possibly leads to suboptimal results, as illustrated in Figure 1. Therefore, it is critical to establish a theoretical framework to answer the question: what is the optimal transfer quantity of samples from each source task needed in training the target task?
In this work, we formulate a sample-based multi-source transfer learning problem as a parameter estimation problem. By employing high-dimensional statistical methods to analyze it, we establish a theoretical framework to determine the optimal transfer quantity for each source task. Specifically, we introduce the expectation of Kullback-Leibler (K-L) divergence between the true distribution of target task samples and the distribution learned from training samples as a measure of generalization error. This measure is then minimized in the asymptotic regime to derive the optimal transfer quantity of each source task. Building on this, we propose a practical algorithm, named Optimal Transfer Quantities for Multi-Source Transfer learning (OTQMS), to implement our theoretical results for training multi-source transfer learning models. Notably, OTQMS is data-efficient and compatible with various model architectures, including Vision Transformer (ViT) and Low-Rank Adaptation (LoRA). It also demonstrates advantages in task generality and shot generality (as illustrated in Table 1), since we establish theoretical framework without restricting it to a specific target task type or limiting the range of target sample quantity. In summary, our main contributions are as follows:
  • a)
    We use high-dimensional statistics to analyze the parameter estimation problem formulated by the sample-based multi-source transfer learning problem. Based on this, a novel theoretical framework that optimizes transfer quantities is introduced.
  • b)
    Based on the framework, we propose OTQMS, an architecture-agnostic and data-efficient algorithm for target model training in multi-source transfer learning. In particular, we propose a dynamic strategy in OTQMS to alleviate the estimation error of transfer quantities caused by the scarcity of target task samples.
  • c)
    Experimental studies on few-shot multi-source transfer learning tasks, two real-world datasets and various model architectures demonstrate that OTQMS achieves significant improvements in both accuracy and data efficiency. In terms of accuracy, OTQMS outperforms state-of-the-art approaches by an average of $1.5\%$ on DomainNet and $1.0\%$ on Office-Home. In terms of data efficiency, OTQMS reduces the average training time by $35.19\%$ and the average sample usage by $47.85\%$ on DomainNet. Furthermore, extensive supplementary experiments demonstrate that OTQMS can facilitate both full model training and parameter-efficient training, and OTQMS is also applicable to multi-task learning tasks.
@@
MCW [12] Supervised
Leep [17] Supervised
Tong [25] Supervised
DATE [9] Unsupervised
H-ensemble [31] Supervised
@@ 2 Related Work
This work is related to two main lines of research. The first is transfer learning theory, where we propose a measure based on K-L divergence as a novel measure of generalization error, different from previously adopted measures  [10, 2, 25, 1, 31]. The second is multi-source transfer learning, where most existing studies either utilize all samples from all source tasks or adopt task-level selection strategies [8, 22, 25, 31]. In contrast, our framework explicitly optimizes the transfer quantity from each individual source task. Other closely related work is discussed further in Appendix B.
@@ 3 Problem Formulation
Consider the transfer learning setting with one target task $\mathcal{T}$}, and $K$ source tasks $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$. The target task $\mathcal{T}$ is not restricted to a specific downstream task category. Generally, we formulate it as a parameter estimation problem under a distribution model $P_{X;{\underline{\theta}}}$. For example, when $\mathcal{T}$} is a supervised classification task, $P_{X;{\underline{\theta}}}$ corresponds to the joint distribution model of input features $Z$ and output labels $Y$, i.e., $X=(Z,Y)$. Our objective is to estimate the true value of $\underline{\theta}$, which corresponds to optimizing the neural network parameters for target task $\mathcal{T}$}. Here, $\theta$ denotes 1-dimensional parameter, and $\underline{\theta}$ denotes high-dimensional parameter. Furthermore, we assume that the source tasks and the target task follow the same parametric model and share the same input space $\mathcal{X}$. Without loss of generality, we assume $\mathcal{X}$ to be discrete primarily for clarity of writing and to avoid redundancy, and our results can be readily extended to continuous spaces. The target task $\mathcal{T}$} has $N_{0}$ training samples $X^{N_{0}}=\{x_{1},\dots,x_{N_{0}}\}$ i.i.d. generated from some underlying joint distribution $P_{X;{\underline{\theta}}_{0}}$, where the parameter ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$. Similarly, the source task $\mathcal{S}_{i}$ has $N_{i}$ training samples $X^{N_{i}}=\{x^{i}_{1},\dots,x^{i}_{N_{i}}\}$ i.i.d. generated from some underlying joint distribution $P_{X;{\underline{\theta}}_{i}}$, where $i\in[1,K]$, and the parameter ${\underline{\theta}}_{i}\in\mathbb{R}^{d}$. In this work, we use the Maximum Likelihood Estimator (MLE) to estimate the true target task parameter $\theta_{0}$. Moreover, the following lemma characterizes the asymptotic behavior of MLE.
Lemma 1.
(Asymptotic Normality of the MLE) [29] When we use MLE only based on target task samples to estimate $\theta_{0}$, i.e.
$\hat{{\underline{\theta}}}_{MLE}=\operatorname*{arg\,max}_{{\underline{\theta}}}{\frac{1}{N_{0}}\sum_{x\in X^{N_{0}}}}\log P_{X;{\underline{\theta}}}(x),$ (1)
under appropriate regularity conditions, the following holds:
$\sqrt{N_{0}}\left(\hat{{\underline{\theta}}}_{MLE}-\underline{\theta}_{0}\right)\xrightarrow{d}\mathcal{N}\left(0,J({\underline{\theta}}_{0})^{-1}\right),$ (2)
where “1{-1}” denotes the matrix inverse and the $J({\underline{\theta}})$ is the Fisher information matrix defined as:
@@
$J({\underline{\theta}})^{d\times d}$ $=\mathbb{E}\left[\left(\frac{\partial}{\partial{\underline{\theta}}}\log P_{X;{\underline{\theta}}}\right)\left(\frac{\partial}{\partial{\underline{\theta}}}\log P_{X;{\underline{\theta}}}\right)^{T}\right].$ (3)
@@
When we transfer $n_{1},\dots,n_{K}$ samples from the $\left\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\right\}$, where $n_{i}\in\left[0,N_{i}\right]$, we denote these training sample sequences as $X^{n_{1}},X^{n_{2}},\dots,X^{n_{K}}$ . Then, the MLE for multi-source transfer learning, is the parameter value that maximizes the empirical mean of the likelihood of samples from source tasks and target task, i.e.,
$\hat{{\underline{\theta}}}=\operatorname*{arg\,max}_{{\underline{\theta}}}\frac{1}{N_{0}+\sum_{i=1}^{K}n_{i}}\left({\sum_{x\in X^{N_{0}}}}\log P_{X;{\underline{\theta}}}(x)+{\sum_{i=1}^{K}\sum\limits_{x\in X^{n_{i}}}}\log P_{X;{\underline{\theta}}}(x)\right).$ (4)
@@
In this work, our goal is to derive the optimal transfer quantities $n_{1}^{*},\dots,n_{K}^{*}$ of source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ to minimize certain divergence measure $\mathcal{D}iv$ between the true distribution of target task $P_{X;{\underline{\theta}}_{0}}$ and the distribution $P_{X;\hat{{\underline{\theta}}}}$ learned from training samples, i.e.,
[…]
@@ 4 Theoretical Analysis and Algorithm
In this section, we will first introduce a new K-L divergence based measure for the optimization problem in (5). Then, we will analyze it based on high-dimensional statistics to derive the optimal transfer quantities for both single-source and multi-source scenarios. Finally, we will develop a practical algorithm based on the theoretical framework.
Definition 2.
(The K-L divergence) [4] The K-L divergence $D\left(P\middle\|Q\right)$ measures the difference between two probability distributions $P(X)$ and $Q(X)$ over the same probability space. It is defined as:
@@
In this work, we apply the expectation of the K-L divergence between the true distribution model of the target task and the distribution model learned from training samples to measure the generalization error.
Finally, our generalization error measure is defined as:
[…]
@@ 4.1 Single-Source Transfer Learning
The direct computation of the proposed K-L divergence based measure is challenging. Fortunately, we can show that the proposed measure is computable in the asymptotic regime using high-dimensional statistical analysis. To be specific, we can prove that the proposed measure directly depends on the Mean Squared Error (MSE) in the asymptotic regime, and the MSE can be calculated by an extension of Lemma 1. Therefore, we perform an asymptotic analysis of this measure.
In this section, we begin by presenting the theoretical results in Lemma 3 and Theorem 4 where the parameter is 1-dimensional, and subsequently extend them to the high-dimensional parameter setting in Proposition 5 and 6. To begin, we consider the setting with a target task $\mathcal{T}$} with $N_{0}$ samples generated from a model with 1-dimensional parameter.
Lemma 3.
(proved in Appendix C.1) Given a target task $\mathcal{T}$} with $N_{0}$ i.i.d. samples generated from a 1-dimensional underlying model $P_{X;\theta_{0}}$, where $\theta_{0}\in\mathbb{R}$, and denoting $\hat{\theta}$ as the MLE (1) based on the $N_{0}$ samples, then the proposed measure (6) can be expressed as:
$\mathbb{E}\left[D\left(P_{X;\theta_{0}}\middle\|P_{X;\hat{\theta}}\right)\right]=\frac{1}{2N_{0}}+o\left(\frac{1}{N_{0}}\right).$ (7)
The result of Lemma 3 demonstrates that when there is only one target task without any source task, the proposed measure is inversely proportional to the number of training samples. Next, we consider the transfer learning scenario where we have one target task $\mathcal{T}$} with $N_{0}$ training samples and one source task $\mathcal{S}_{1}$ with $N_{1}$ training samples. In this case, we aim to determine the optimal transfer quantity $n_{1}^{*}\in[1,N_{1}]$. To facilitate our mathematical derivations, we assume $N_{0}$ and $N_{1}$ are asymptotically comparable, and the distance between the parameters of the target task and source task is sufficiently small (i.e., $|\theta_{0}-\theta_{1}|=O(\frac{1}{\sqrt{N_{0}}})$). Considering the similarity of low-level features among tasks of the same type, this assumption is made without loss of generality and supported by related studies [20]. Furthermore, as demonstrated in subsequent analysis, our conclusions remain valid even in extreme cases where the distance between parameters is large.
Theorem 4.
(proved in Appendix C.2) Given a target task $\mathcal{T}$} with $N_{0}$ i.i.d. samples generated from a 1-dimensional underlying model $P_{X;\theta_{0}}$, and a source task $\mathcal{S}_{1}$ with $N_{1}$ i.i.d. samples generated from a 1-dimensional underlying model $P_{X;\theta_{1}}$, where $\theta_{0},\theta_{1}\in\mathbb{R}$ and $|\theta_{0}-\theta_{1}|=O(\frac{1}{\sqrt{N_{0}}})$, $\hat{\theta}$ is denoted as the MLE (4) based on the $N_{0}$ samples from $\mathcal{T}$} and $n_{1}$ samples from $\mathcal{S}_{1}$, where $n_{1}\in[0,N_{1}]$, then the proposed measure (6) can be expressed as:
$\frac{1}{2}\bigg(\underbrace{\frac{1}{N_{0}+n_{1}}}_{\textnormal{variance~ term}}+\underbrace{\frac{n_{1}^{2}}{(N_{0}+n_{1})^{2}}t}_{\textnormal{bias~term}}\bigg)+o\left(\frac{1}{N_{0}+n_{1}}\right),$ (8)
@@
From (8), we observe that the proposed measure decreases as $N_{0}$ increases, which aligns with our intuition that utilizing all available target samples is beneficial. In addition, the trend of (8) with respect to $n_{1}$ is more complex. We plot (8) as a function of $n_{1}$ under two different regimes, determined by the value of $N_{0}\cdot t$, as shown in Figure 2. Our goal is to explore the optimal value of $n_{1}^{*}$ to minimize (8).
Figure 2: The function curve figures of (8) under different regimes determined by the value of $N_{0}\cdot t$ (blue). The vertical axis denotes the value of proposed measure (8), while the horizontal axis denotes the variable $n_{1}$.
  • Case 1 ($N_{0}\cdot t\leq 0.5$): The proposed measure monotonically decreases as $n_{1}$ increases. Obviously, the optimal point is $n_{1}^{*}=N_{1}$. This indicates that when the source task and the target task are highly similar, i.e., when $t$ is small, an increase in the transfer quantity will positively impact the results.
  • Case 2 ($N_{0}\cdot t>0.5$): The proposed measure first decreases and then increases as $n_{1}$ increases. It attaining its minimum at $n_{1}^{\prime}=\frac{N_{0}}{2N_{0}t-1}$. It should be noted that when $t$ is large enough, the point $n_{1}^{\prime}$ approaches 0. This aligns with the intuition that when the discrepancy between the source and target tasks is substantial, avoiding transfer yields better results. Furthermore, if $n_{1}^{\prime}$ exceeds $N_{1}$, we should utilize all $N_{1}$ samples, so $n_{1}^{*}=\min\left(N_{1},\frac{N_{0}}{2N_{0}t-1}\right)$.
@@
Proposition 5.
(proved in Appendix C.3) In the case where the parameter dimension is d, i.e., ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$, with all other settings remaining the same as the Lemma 3, the proposed measure (6) can be expressed as
@@
Proposition 6.
(proved in Appendix C.4) In the case where the parameter dimension is d, i.e., ${\underline{\theta}}_{0},{\underline{\theta}}_{1}\in\mathbb{R}^{d}$, with all other settings remaining the same as the Theorem 4, the proposed measure (6) can be expressed as